OpenStudy (anonymous):

What are the roots of 4x2 + 176 = 0?

OpenStudy (anonymous):

\[4x ^{2}=-176,x ^{2}=\frac{ -176 }{ 4 }=44\times-1=44\iota ^{2},x=\pm2\sqrt{11}\iota \]